Enthalpy of N2O4--> + 20
1
Expert's answer
2016-12-04T03:37:09-0500
1) N2(g) + 202(g) → 2NO2(g) ∆H =+66.36 kJ

2) N2(g) + 202(g) → N2O4(g) ∆H = +9.16 kJ


N2O4(g) → N2(g) + 202(g) ∆H = -9.16 kJ
N2(g) + 202(g) → 2NO2(g) ∆H = +66.36 kJ

N2O4(g) → 2NO2(g) ∆H = +57.2 kJ

Enthalpy of N2O4 is ∆H = +57.2 kJ